OpenStudy (anonymous):

What is the Celsius temperature of 1 mole of a gas that has an average kinetic energy of 4,290 joules?

OpenStudy (mathstudent55):

\(\large KE_{avg} = \dfrac{3}{2}RnT\) \(R = universal~gas ~constant = 8.31 \dfrac{J}{K \cdot mol} \) \(n = number~of~moles\) \(T = temperature~in~Kelvin\) Use the formula to find the Kelvin temperature, and then convert to Celsius.

OpenStudy (anonymous):

thanks, i was forgetting to convert to kelvin

OpenStudy (mathstudent55):

You're welcome.